Circular No. 6104

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

X-RAY NOVA 1993 IN OPHIUCHUS B. A. Harmon, S. N. Zhang, W. S. Paciesas, C. A. Wilson, and G. J.

Fishman report for the Compton Observatory BATSE Team: “This object (GRS 1716–249 = GRO J1719–24) has gradually brightened since the reported detection in early September by GRANAT/Sigma (IAUC 6083) to ∼ 300 mCrab in the band 20–100 keV, as of Nov. 7. The spectrum hardened steadily from early September to ∼ Sept. 30, with power-law spectral index (20–100 keV) increasing from –2.5 to –1.8. In October, the spectrum was marginally softer, with index –2.2 to –2.4. The spectrum shows a steepening above 100 keV and extends to at least 200 keV. The light curve, spectrum, and spectral evolution are similar to a long-lived secondary maximum detected in June-September 1991 for Nova Mus 1991 (Paciesas et al. 1994, AIP Conf. Proc. 308, 115); however, there is no clear peak in the light curve of the x-ray nova in Ophiuchus at this time. We note that this outburst follows the initial outburst in September 1993 (IAUC 5874) by ∼ 11 months.”

OJ 287 M. Fiorucci and G. Tosti, Department of Physics, University of Pe-

rugia, report that their further photometry confirms that this object is at its brightest level since 1984 (cf. IAUC 6095): Oct. 13.15 UT, V = 14.25; 14.11, 14.29; 15.16, 14.56; 16.12, 14.35; 18.17, 14.54; 19.13, 14.44; 25.13, 14.39; 31.14, 14.33; Nov. 5.15, 14.63; 12.08, 14.11. They also note the presence of oscillations having amplitudes of ∼ 0.5 mag in V and durations of few days. Additional photometry communicated by T. Pursimo, P. Heinämäki, H. Lehto, A. Sillanpää, and L. Takalo (Tuorla Observatory); and J. A. de Diego, M. Kidger, and I. Vidal (Instituto de Astrofisica de Canarias), made partly with Indiana University’s Automatic RoboScope telescope: Nov. 6.01 UT, V = 14.08; 6.44, 14.14; 8.47, 14.21; 10.45, V = 13.99, R = 13.60, I = 13.02 (P. Boltwood, Canada); 11.08, 14.02, –, –; 11.26, 13.99, 13.56, 13.01; 11.39, 13.97, –, –.

3C 66A Additional V photometry (cf. IAUC 6095) by Fiorucci and Tosti: Oct.

13.10 UT, 14.28; 13.98, 14.48; 14.02, 14.50; 15.09, 14.47; 16.05, 14.51; 18.12, 14.22; 19.00, 14.20; 22.93, 14.16; 25.06, 14.34; 25.90, 14.37; 27.84, 14.02; 30.86, 14.11; Nov. 3.06, 13.81; 4.97, 14.06; 12.06, 13.89.

1994 November 14 Daniel W. E. Green